A bus with a mass of 809 kg is moving at a constant speed of +5.1 m/s in a straight line. It is brought to rest in 22.0 seconds. The constant force is acting to stop the car is?


1
Expert's answer
2021-11-18T15:09:32-0500

"F=ma"


"v=v_0+at\to a=(v-v_0)/t=(0-5.1)/22=0.23\ (m/s)"


"F=809\cdot0.23=187.5\ (N)" . Answer